OpenStudy (anonymous):

A right triangle has base (x - 6) units and height (6x - 12) units.

OpenStudy (anonymous):

Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. Part B: What is the area of the triangle? Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials.

OpenStudy (anonymous):

Only answer A through messages please :D

OpenStudy (alexandervonhumboldt2):

area: (6x-12)*(x-6)/2

OpenStudy (anonymous):

mhm....

OpenStudy (anonymous):

I can see you got that by multiplying (x - 6) by (6x - 12) right?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

ok, what next?

OpenStudy (alexandervonhumboldt2):

and then divde by 2

OpenStudy (tylerd):

1/2 base times height = right triangle area

OpenStudy (anonymous):

yeah.... oh thanks :D

OpenStudy (alexandervonhumboldt2):

1/2*b*h is the area of any triangle

OpenStudy (anonymous):

x=3x^2−24x+36 right?

OpenStudy (anonymous):

(refresh page)

OpenStudy (alexandervonhumboldt2):

hypotenuse^2=(x-6)^2+(6x-12)^2

OpenStudy (anonymous):

what is the hypotenuse?

OpenStudy (tylerd):

hero formula is the area of any triangle

OpenStudy (anonymous):

what is the hypotenuse in your case AlexandervonHumboldt2?

OpenStudy (alexandervonhumboldt2):

hypotenuse^2=(x-6)^2+(6x-12)^2

OpenStudy (alexandervonhumboldt2):

hypotenuse=sqrt((x-6)^2+(6x-12)^2)

OpenStudy (anonymous):

after trying multiple times, I don't believe it's possible to solve that without knowing what x is @AlexandervonHumboldt2 !

OpenStudy (alexandervonhumboldt2):

it cannot be known

OpenStudy (alexandervonhumboldt2):

the answer is with variable

OpenStudy (anonymous):

ah ok.... so mind helping me out?

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so what do we start by doing?

OpenStudy (alexandervonhumboldt2):

what you need help with?

OpenStudy (anonymous):

solving for the hypotenuse :|

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

so what do we start by doing?

OpenStudy (alexandervonhumboldt2):

so in part A we need the sqruare of hypotenuse so...

OpenStudy (alexandervonhumboldt2):

hypotenuse^2=(x-6)^2+(6x-12)^2

OpenStudy (anonymous):

ok...

OpenStudy (alexandervonhumboldt2):

x^2-12x+36+36x^2-144x+144

OpenStudy (anonymous):

37x^2+−156x+180

OpenStudy (alexandervonhumboldt2):

37x^2-156x+180

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

is that the hyp or still more to do?

OpenStudy (alexandervonhumboldt2):

square of hypotenuse is 37x^2-156x+180

OpenStudy (anonymous):

ah ok... so A?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

off to b?

OpenStudy (alexandervonhumboldt2):

area: (6x-12)*(x-6)/2

OpenStudy (anonymous):

haha ok then, off to c?

OpenStudy (alexandervonhumboldt2):

(6x-12)*(x-6)/2=(6x^2-48x+72)/2=3x^2-24x+36

OpenStudy (anonymous):

ok wow that is a lot of calculating :D

OpenStudy (alexandervonhumboldt2):

not a lot had you fanned me?

OpenStudy (alexandervonhumboldt2):

thx now

OpenStudy (anonymous):

Haha yw :D

OpenStudy (anonymous):

Can you through messages help me explain the closure property?

OpenStudy (alexandervonhumboldt2):

probably no because i dont know closure property

OpenStudy (alexandervonhumboldt2):

sorry

OpenStudy (anonymous):

no that's ok, but do you know anyone who does?

OpenStudy (alexandervonhumboldt2):

@ganeshie8

OpenStudy (alexandervonhumboldt2):

@Directrix

OpenStudy (alexandervonhumboldt2):

@TylerD

OpenStudy (anonymous):

thanks :D

OpenStudy (alexandervonhumboldt2):

i mesagged ganishe lets see

OpenStudy (anonymous):

ok :)

OpenStudy (tylerd):

?

OpenStudy (anonymous):

hi :D

OpenStudy (tylerd):

oh closure property

OpenStudy (anonymous):

yeah :D

OpenStudy (alexandervonhumboldt2):

i found explanation: A set has closure under an operation if performance of that operation on members of the set always produces a member of the same set; in this case we also say that the set is closed under the operation.

OpenStudy (anonymous):

hmm... somewhat complicated, but I understand some parts...

OpenStudy (alexandervonhumboldt2):

A set that is closed under an operation or collection of operations is said to satisfy a closure property.

OpenStudy (anonymous):

how would we explain that using b?

OpenStudy (alexandervonhumboldt2):

probably as b has no limits

OpenStudy (anonymous):

oh really... why?

OpenStudy (tylerd):

no clue

OpenStudy (alexandervonhumboldt2):

well because in ^2 everything is poitie lenght cannot be negative

OpenStudy (anonymous):

ah ok...

OpenStudy (anonymous):

thanks a lot.... :D

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (alexandervonhumboldt2):

i had becomed green today

OpenStudy (tylerd):

its probably something dumb like hieght/2 times base/2 = blah

OpenStudy (alexandervonhumboldt2):

haha

OpenStudy (anonymous):

oh and congrats @AlexandervonHumboldt2 ... I can see, you're green not lime :D

OpenStudy (alexandervonhumboldt2):

well yeah thx

OpenStudy (alexandervonhumboldt2):

i posted that in study group OpensTudy feedback

OpenStudy (anonymous):

:D green and mean...

OpenStudy (anonymous):

:D green and mean...